# Do Graph Neural Networks Dream of Landau Damping? Insights from Kinetic Simulations of a Plasma Sheet Model

### v2

@@ 6.5 Landau Damping
While fast sheets are able to excite an electrostatic wake in their rear, the resulting electrostatic wake can also accelerate sheets moving slightly slower than its phase velocity [1, 23]. Electrostatic modes are therefore self-consistently generated by particles moving close to its phase velocity, while being damped by particles moving slightly slower. However, since a plasma in thermal equilibrium follows a Maxwellian distribution in velocity space, there exist on average more particles moving faster than the wave, than those moving slower. Therefore, on average, the modes will be damped. This mechanism is known as Landau damping [1] and is an inherently collisionless kinetic process that the sheet model has been shown to recover [23].
For a given mode $m$, with a wavelength $\lambda_{m}=2L/m$ and wave vector $k_{m}=2\pi/\lambda_{m}$, we can compute its wave frequency and damping time by finding numerically the roots of the dispersion relation for $k=k_{m}$ [39]:
[…]
where $\hat{f}_{0}$ corresponds to the distribution function in velocity space. The solution will have both a real and imaginary part. The real part corresponds to the wave angular frequency, while the imaginary part corresponds to the inverse of the damping time.
To reproduce the expected damping behavior we follow a similar procedure to that of Dawson [23]. We produce 50 simulations, each with a duration of $t_{max}=500~{}\omega_{p}^{-1}$, of thermal plasmas consisting of $10^{3}$ sheets for $\lambda_{D}=5\,\delta$ and reflecting boundaries. For each simulation time step we compute the mode “amplitude” $A_{m}$ and its rate of change $\dot{A}_{m}$ using the cross-correlation:
@@
In figure 14 we showcase the results obtained for several modes. It is possible to see that, although for some trajectories the mode is still growing, on average it decreases according to the expected damping time. This demonstrates that the GNS is capable of correctly modeling Landau damping, an inherently kinetic mechanism associated with the collective collisionless dynamics of a plasma.
Figure 14: Damping of different modes. The initial mode amplitudes are normalized to $R=0.3\,\delta$. Individual trajectories length is $\Delta t_{traj}=50~{}\omega_{p}^{-1}$. Damping times are respectively $\tau_{28}=10.25~{}\omega_{p}^{-1}$, $\tau_{30}=7.97~{}\omega_{p}^{-1}$, and $\tau_{32}=6.41~{}\omega_{p}^{-1}$. The agreement between the theoretical curves and the average mode trajectories demonstrates that the GNS is capable of correctly modeling Landau damping, an inherently kinetic mechanism.
@@ 6.6 Two-stream Instability
Finally, we provide in H further comparisons for higher order modes (lower $v_{0_{m}}$) and larger simulation steps ($\Delta t=10^{-2}~{}\omega_{p}^{-1}$), as well as results for the remaining GNNs trained with different seeds (with worse energy conservation capabilities at high $v_{th}$). These results further support the claim that the GNS is consistently able to correctly model the overall dynamics of the instability.
@@ 7 Conclusions
It is also important to note that, despite being very simple, the sheet model provides a powerful framework to explore fundamental plasma processes, as demonstrated in this paper, including the relevance of collisional processes. Furthermore, direct extensions and/or similar models exist, which allow one to study multi-species scenarios [24, 28], electron-neutral collisions [41], and electron-ion 3D scattering events [42]. Therefore, the demonstration that the GNS can accurately provide a surrogate for the sheet model can open new directions for the exploration of the fundamentals of statistical mechanics of plasmas, and further extensions of the framework can be developed to integrate multi-species and stochastic processes.
Lastly, this work suggests that it would be possible to accurately model the standard PIC loop. However, we do not believe that this would result in meaningful computational gains since the standard PIC loop implemented in modern architectures is extremely optimized and does not suffer from the issues that are present in the sheet model scenario (e.g. time-consuming serial routines). Instead, the use of graph neural networks can be relevant to model multi-scale plasma dynamics or extra physics modules. This will be explored in future publications.
@@
The code used for this project is made available through a git repository: (to be disclosed). The training and test data, as well as a set of the final model weights are available in a zenodo repository: (to be disclosed)
@@
The authors would like to thank P. J. Bilbao, T. Grismayer, A. Joglekar, and E. P. Alves for helpful discussions. This work was supported by the FCT - Fundação para a Ciência e Tecnologia, I.P. under the Project No 2022.02230.PTDC (X-MASER) and PhD Fellowship Grant 2022.13261.BD and has received funding from the European Union’s H2020 programme through the project IMPULSE (grant agreement No 871161). The graphics processing units (GPUs) used in this work were donated by NVIDIA Corporation.
@@ A.1 Synchronous Algorithm (SM)
where $-v^{max}_{-}\leq 0$ is the maximum negative velocity across all sheets, whose value is updated every $\omega_{p}^{-1}$. However, we found that this approximation is not sufficient for non-thermal initial conditions with very large sheet velocities and displacements from equilibrium (e.g. the two-stream instability simulations shown in Section 6.6 and H) since it underestimates the real $\Delta x^{t}_{max_{i}}$ value. As a consequence, some crossings are not detected, which causes problems when trying to sort sheets after the crossing corrections (i.e. multiple sheets would occupy the same array index). For this reason, we instead derive a more accurate upper limit from (14) that is:
@@ Appendix G Model Benchmark – Extra
Appendix H Two-stream Instability – Extra

### v3

@@ 6.6 Landau damping
While fast sheets are able to excite an electrostatic wake in their rear, the resulting electrostatic wake can also accelerate sheets moving slightly slower than its phase velocity [1, 37]. Electrostatic modes are therefore self-consistently generated by particles moving close to its phase velocity, while being damped by particles moving slightly slower. However, since a plasma in thermal equilibrium follows a Maxwellian distribution in velocity space, there exist on average more particles moving faster than the wave, than those moving slower. Therefore, on average, the modes will be damped. More specifically, for a given mode $m$, with a wavelength $\lambda_{m}=2L/m$ and wave vector $k_{m}=2\pi/\lambda_{m}$ we can compute its wave frequency and damping time by finding numerically the roots of the dispersion relation (11). This damping mechanism is known as Landau damping [1, 55] and is an inherently collisionless kinetic process that the sheet model has been shown to recover [37, 41].
To reproduce the expected damping behavior we follow initially a similar procedure to that of Dawson [37]. We produce 100 simulations, each with a duration of $t_{max}=500~{}\omega_{p}^{-1}$, of thermal plasmas consisting of $10^{4}$ sheets for $\lambda_{D}=5\,\delta$ and reflecting boundaries. For each simulation time step we compute the mode “amplitude” $A_{m}$ and its rate of change $\dot{A}_{m}$ using the cross-correlation:
@@
In figure 17 we showcase examples of the results obtained for several modes. It is possible to see that, although for some trajectories the mode is still growing, on average it decreases according to the expected damping time. This demonstrates that the GNS is capable of correctly modeling Landau damping, an inherently kinetic mechanism associated with the collective collisionless dynamics of a plasma.
Figure 17: Damping of different modes. The initial mode amplitudes are normalized to $R=0.15\,\delta$. Individual trajectories length is $\Delta t_{traj}=50~{}\omega_{p}^{-1}$. The agreement between the theoretical curves and the average mode trajectories demonstrates that the GNS is capable of correctly modeling Landau damping, an inherently kinetic mechanism.
To further support this claim, we perform a scan over different modes $m\in[180,448]$ ($k\lambda_{D}\in[0.28,0.70]$) and initial mode amplitudes $A^{0}_{m}\in[0.08,0.20]~{}\delta$ using a resolution of $\Delta m=2$ and $\Delta A^{0}_{m}=0.005~{}\delta$ ($135\times 25=3375$ data points). For each ($m$, $A^{0}_{m}$) pair, we compute the average of the trajectories in phase-space (figure 17). We then proceed to estimate the damping rate, $-\mathrm{Im}(\omega)$, by obtaining the slope of the line that best fits the peaks of $\log|A^{t}_{m}|$. The mode angular frequency, $\mathrm{Re}(\omega)$, is estimated as $\pi/\Delta t^{avg}_{peaks}$, where $\Delta t^{avg}_{peaks}$ is the average interval between consecutive peaks of $\log|A^{t}_{m}|$ (more details in H).
The Landau damping rates are shown in figure 18, and the corresponding angular frequency (the real part of $\omega$) in figure 19. The average values obtained by the GNS are in good agreement with the theory, and both the damping rate and angular frequency are (as expected) independent of the initial mode amplitude. The higher variations observed for larger values of $k\lambda_{D}$ are due to having less periods for the estimation of $\omega$ since these correspond to strongly damped modes (for $k\lambda_{D}>0.6$ the phase velocity $v_{ph}\leq 2.6~{}v_{th}$, already in the plasma bulk).
Figure 18: Comparison between the theoretical mode damping rate, $-\mathrm{Im}(\omega)$, and the one measured using the GNS. Results are shown for different mode angular wavenumbers and initial mode amplitudes (no interpolation is performed), as well as the average damping rate across initial mode amplitudes (standard deviation in lighter color). It is observed that the GNS correctly approximates the theoretical damping rate for a broad range of modes.
Figure 19: Comparison between the theoretical mode angular frequency, $\mathrm{Re}(\omega)$, and the one measured using the GNS. Similarly to figure 18, we present results for different mode angular wavenumbers and initial mode amplitudes, as well as the average angular frequency across initial mode amplitudes. Although there is a slight underestimation of the angular frequency, the GNS is capable of approximating the theoretical prediction across a broad range of modes.
@@ 6.7 Two-stream instability
Finally, we provide in I further comparisons for higher order modes (lower $v_{0_{m}}$) and smaller simulation steps ($\Delta t=10^{-2}~{}\omega_{p}^{-1}$), as well as results for the remaining GNNs trained with different seeds (with worse energy conservation capabilities at high $v_{th}$). These results further support the claim that the GNS is consistently able to correctly model the overall dynamics of the instability.
@@ 7 Conclusions
It is also important to note that, despite being very simple, the sheet model provides a powerful framework to explore fundamental plasma processes, as demonstrated in this paper, including the relevance of collisional processes. Not only is it a gridless code (thus making fewer approximations than PIC codes) but also provides an exact description of one-dimensional non-relativistic electrostatic plasmas with arbitrary degrees of collisionality. Furthermore, direct extensions and/or similar models exist, which allow one to study multi-species scenarios [38, 42], electron-neutral collisions [57], and electron-ion 3D scattering events [58]. Therefore, the demonstration that the GNS can accurately provide a fast surrogate for the sheet model can open new directions for the exploration of the fundamentals of statistical mechanics of plasmas, and further extensions of the framework can be developed to integrate multi-species and stochastic processes.
Additionally, this work suggests that it would be possible to design and train a GNS that accurately models the standard PIC loop. This could be done by extending the developed framework to include grid information in both the graph representation and GNN architecture [23]. However, the standard PIC loop implemented in modern architectures is extremely optimized and does not suffer from the issues that are present in the sheet model scenario (e.g. time-consuming serial routines). In fact, the PIC loop itself can already be seen from a graph computation perspective: field interpolation and particle push can be seen as message-passing steps from the grid to the particles; current/charge deposition as a message-passing step from the particles to the grid; and the field solver as a message-passing step between neighboring cells. Note that, frequently, most of these operations are linear. Therefore, simply replacing them by a learnable function is counter-productive and should lead to an expected increase in run-time.
The question is then, where might a GNN-based approach introduce some gain compared with traditional PIC? In our view, two main possibilities arise: a) GNNs could be used to relax the time constraints imposed by the CFL condition [3]; b) GNNs could introduce extra physics that the standard PIC loop fails to model, which usually require additional modules (e.g. collisions – thus having a model that intrinsically addresses collisions is important on this roadmap). Approaches following the first route should ensure that causality is still verified (similarly to what we have shown for the sheet model), this will result necessarily in increased graph connectivity (e.g. particles would require information from neighboring grid cells) and will probably require the exploration of multi-scale approaches [28, 27] or hybrid representations [16] to be competitive. The second route will result in an increased run-time compared to the standard PIC loop but might be competitive against alternative approaches. For example, a standard 1D1V PIC simulation does not correctly capture collisional dynamics. Again, the sheet model, and consequently the GNS framework presented in this work, can do so for arbitrary levels of collisionality, thus providing a valuable source of comparison.
Similarly to what we have done for the sheet model, it would also be important to explore the trade-off between how much is learned by the GNN, and how much prior knowledge about kinetic simulations is embedded into the GNS. As one moves to higher-dimensional and more complex scenarios, this will become even so more important to maximize generability and possibly reduce the run-time. We will further explore these topics in future publications.
@@
The code used for this project is openly available in a git repository: https://github.com/diogodcarvalho/gns-sheet-model. The training and test data, as well as a set of the final model weights, are available in a zenodo repository: http://doig.org/10.5281/zenodo.10440186.
@@
The authors would like to thank P. J. Bilbao, B. Malaca, T. Grismayer, A. Joglekar, E. P. Alves, and V. Decyk for helpful discussions and the anonymous referees for their valuable comments which helped improve the quality of this manuscript. This work was supported by the FCT - Fundação para a Ciência e Tecnologia, I.P. under the Project No 2022.02230.PTDC (X-MASER) and PhD Fellowship Grant 2022.13261.BD and has received funding from the European Union’s H2020 programme through the project IMPULSE (grant agreement No 871161). The graphics processing units (GPUs) used in this work were donated by NVIDIA Corporation.
@@ A.1 Synchronous Algorithm (SM)
where $-v^{max}_{-}\leq 0$ is the maximum negative velocity across all sheets, whose value is updated every $\omega_{p}^{-1}$. However, we found that this approximation is not sufficient for non-thermal initial conditions with very large sheet velocities and displacements from equilibrium (e.g. the two-stream instability simulations shown in Section 6.7 and I) since it underestimates the real $\Delta x^{t}_{max_{i}}$ value. As a consequence, some crossings are not detected, which causes problems when trying to sort sheets after the crossing corrections (i.e. multiple sheets would occupy the same array index). For this reason, we instead derive a more accurate upper limit from (14) that is:
@@ Appendix G Model benchmark – Extra
Appendix H Landau damping – Extra
The range used for the initial mode amplitudes $A_{m}^{0}\in[0.08,0.2]~{}\delta$ in Section 6.6 was chosen so that: a) the minimum initial amplitude was above the usual “noise” level, i.e. the values at which on average most nodes stop decaying; b) the maximum value was not too high to guarantee we had enough statistics (number of trajectories) to obtain a good estimate of the average trajectory across all modes. To further illustrate the reasoning behind the chosen range we provide in figure 38 the amplitude of several modes over time for one of the simulations produced in Section 6.6.
Figure 38: Examples of mode amplitude evolution for one of the simulations used in Section 6.6. The black dashed lines delimit the region used in the Landau Damping diagnostic results presented in figure 18 and figure 19. This region is chosen such that we start tracking the mode evolution above the typical “noise” level while guaranteeing that enough statistics exist to compute a meaningful average trajectory.
The range of the modes used, $k\lambda_{D}\in[0.28,0.7]$, was selected such that enough particles exist to resolve the distribution function near the mode phase velocity (for $k\lambda_{D}=0.28$, $v_{ph}\simeq 4v_{th}$ meaning only $\simeq 100$ particles have $v>v_{ph}$) and that the mode does not decay too fast (for $k\lambda_{D}=0.7$, $-\mathrm{Im}(\omega)^{-1}\simeq 2.5~{}\omega_{p}^{-1}$). It is therefore expected that, closer to these limits, the dynamics are not as well approximated.
To automatically compute the decay rate and the angular frequency we proceed as follows. We first obtain the average trajectory of the mode in the ($A_{m}$, $\dot{A}_{m}$) phase-space as explained in Section 6.6. The length of this trajectory is chosen so that it corresponds to the maximum of: a) the time it takes for the mode to decay to $1/5$-th of its initial amplitude (using the theoretical estimate of the damping rate); b) 3 mode periods (using the theoretical estimate of the angular frequency). The first condition is the preferred one for slowly damped modes, and the second for fastly damped modes.
After obtaining the average trajectory, we split it into equal regions of size $\Delta t=\pi/\omega_{t}$ where $\omega_{t}$ is the theoretical angular frequency of the mode. For each time interval, we retrieve the maximum value of $|A_{m}^{t}|$ and the corresponding time. This process is illustrated in figure 39. The decay rate is then obtained from the slope of the $\log|A_{m}^{t}|$ maxima curve, and the angular frequency from the average distance between consecutive maxima. For highly damped modes (larger $m$, or equivalently larger $k\lambda_{D}$) this estimate is noisier (e.g. $m=400$ in figure 39), resulting in the larger variations presented in figure 18 and figure 19.
Figure 39: Illustration of the Landau Damping diagnostic used to retrieve the damping rate and angular frequency for figure 18 and figure 19. For each mode $m$ and initial mode amplitude $A^{0}_{m}$ we compute the average mode amplitude over time. The maximum for each half period (shown by a black line) is then automatically calculated. The mode damping rate and the angular frequency are extracted based on these maxima.
Appendix I Two-stream instability – Extra